Rewrite the equation 2x - 8y = 48 2x - 8y = 48 as a function of x.
Please help.
Thank you.

Answer by longjonsilver(2297)

2x - 8y = 48
2x -48 = 8y

or 8y = 2x - 48
--> y = (2/8)x - 48/8
--> y = (x/4) - 6

this is your answer. In function notation, f(x) = (x/4) - 6

jon